OpenStudy (nannie):

show me how to solve this, PLEASE!!! 8/v=9/v-1/3

OpenStudy (anonymous):

Add 1/3 to each side: 8/v + 1/3 = 9/v Get the variables to one side: 1/3 = 9/v - 8/v Simplify: 1/3 = (9-8)/v Cross multiply: v = 3*(9-8) = 3*1 = 3 v = 3